OpenStudy (anonymous):

Find the indicated sum for the arithmetic series. S15for 5 + 9 + 13 + 17 + ...

OpenStudy (campbell_st):

for this equation you need the formula \[S_{n} = \frac{n}{2}[2a +(n - 1)\times d]\] n = number of terms, 15 a = 1st term, 5 and d is the common difference, or what is added to each term to get the next term. Just substitute and then evaluate

OpenStudy (anonymous):

alright I understand thanks!!

OpenStudy (anonymous):

495